# Many-Shot In-Context Learning

### v1

@@ 5.4 NLL May Not Be Predictive of ICL Performance
Figure 14: Negative Log-Likelihood as a function of number of shots We plot negative log-likelihood (NLL) on ground truth test set solutions for GPQA, MATH and GSM8K. For GPQA and MATH, questions for Reinforced ICL and Unsupervised ICL comes from the training splits of those datasets. We study GSM8K in the transfer setting, i.e. questions for Reinforced ICL and Unsupervised ICL come from MATH. Note that the absolute NLL numbers for ICL and Reinforced ICL are not directly comparable to Unsupervised ICL, since they employ different prompt formats.
@@
Prior works (Xiong et al., 2023; Anil et al., 2024; Kaplan et al., 2020) have found that the negative log-likelihood (NLL) for ground-truth test outputs, which corresponds to next-token prediction loss, decreases predictably as the context length increases. We confirm this finding for GPQA, Hendrycks MATH and GSM8K with many-shot in-context learning, and report our results in Figure 14. However, we note that NLL trends are not a strong predictor for downstream task performance. For example, the success rate for both MATH and GPQA with ICL decreases after 125 shots (Figure 7,8), but we do not observe a corresponding increase in the negative log-likelihood in Figure 14.
We also plot negative log-likelihood curves for Reinforced and Unsupervised ICL, and find them to generally have a smaller slope when compared to supervised ICL. Furthermore, we observe that NLL for ICL with ground-truth outputs is much lower compared to model-generated outputs, depsite model-generated outputs often resulting in better performance than ground-truth outputs. These results imply that NLL may not be predictive of downstream ICL performance when using prompts that are out-of-distribution with respect to the test set.
In the GSM8K transfer setting (using MATH problems and solutions to score GSM8K solutions), we see that the change in NLL is close to nil. Once again, these trends do not necessarily reflect the final performance observed in Figure 7, as GSM8K performance continues to increase with more examples in the transfer setting. Finally, while MATH performance for Reinforced and Unsupervised ICL saturates around 25-shot prompts, the NLL continues to decrease with more shots.
@@
Our results lead us to conclude that negative log-likelihood is not a reliable proxy when attempting to predict final performance for problem-solving domains. This makes intuitive sense: for any given problem, there are a large number of potentially correct chain-of-thought paths that the model can take, and calculating the log-likelihood on only one such path may not provide a clear picture for overall model capability. We also explore computing negative log-likelihood on a diverse set of model-generated outputs on MATH, and our findings are presented in Appendix A.1.
6 Conclusion
@@
We investigated many-shot in-context learning. We found significant gains in performance when going from few-shot to many-shot in a wide range of tasks, including translation, summarization, planning, reward modeling, mathematical problem solving, scientific question-answering, and algorithmic reasoning. To overcome the challenges of obtaining a large number of high-quality human-written rationales for many-shot learning, we introduced two new in-context learning regimes: Reinforced ICL and Unsupervised ICL. We found that, for problem-solving domains where human-generated rationales are expensive to obtain, Reinforced and Unsupervised ICL can obtain strong performance when compared to ICL with human data. We concluded with a set of analysis experiments showing that many-shot learning can overcome pre-training biases, allows learning non-natural language tasks typically difficult for LLMs with few-shot learning, and explored negative log-likelihood trends with respect to context length in the many-shot setting.
Limitations
@@
An important limitation of our work is that it uses a single model for all its experiments, the Gemini 1.5 Pro (Gemini Team, 2024). Future work should focus on evaluating the many-shot in-context learning abilities of a wide range of long context models, as they become available. Many-shot performance can likely serve as a valuable metric for evaluating the quality of long-context models, going beyond the needle-in-a-haystack test (Kamradt, 2023) that is currently popular for this purpose.
@@
Another limitation of our work is that we don’t completely understand why performance can sometimes degrades with more examples in the prompt (for example, for MATH). Our analysis found that negative log-likelihood trends are insufficient to explain this degradation, and future work should focus on investigating new research directions to shed light on the matter.
@@ Acknowledgements
We would like to thank Gheorghe Comanici for reviewing an early draft of this work. We are also grateful to Doina Precup, Aviral Kumar, Dale Schuurmans, Ankit Anand, Ross Goroshin, Urvashi Singh, and Daniel Toyoma for helpful discussions.
@@
RA initiated the project, ran majority of the many-shot experiments and analysis, came up with reinforced ICL, on-boarded collaborators, wrote the initial draft. AS contributed initial infra for experiments on MATH and GSM8K, ran BBH experiments, conducted NLL analysis on problem-solving tasks, and wrote several sections of the paper.
LZ contributed results for in-context verifier. BB contributed the planning logistics task. AA helped with GPQA, SC contributed the baseline for parity task and both helped edit the paper. AF and HL provided feedback on an early draft. HL also suggested the unsupervised ICL experiments. Others were involved in project discussions and minor edits to the paper.
@@ Appendix A Appendix
A.1 Negative Log-Likelihood on Model-Generated Data
A.2 Hallucinations with Many-Shot Prompting on Summarization
Here is an example on XSum showing summaries with halluncinated date and time when increasing the number of shots for many-shot learning.
50-shot: 12 September 2016 Last updated at 14:32 GMT One of Russia’s last independent pollsters has been designated a "foreign agent" by the justice ministry.
125-shot: 12 September 2016 Last updated at 14:43 GMT One of Russia’s last independent pollsters has been branded a "foreign agent" by the justice ministry.
250-shot: 2016-09-06 Last updated at 14:33 GMT Russia’s only independent pollster has been branded a "foreign agent" by the justice ministry.
500-shot: 12 September 2016 Last updated at 14:22 GMT One of Russia’s last independent pollsters has been branded a "foreign agent" by the government.
A.3 Additional Results for Reinforced and Unsupervised ICL
Figure A.16: Hendrycks MATH and Minvera Prompt. We find the performance of model-generated rationales with 4-shot Minerva prompt is generally better or comparable to the ones generated by 4-shot InnerMono prompt. Furthermore, another iteration of Reinforced ICL – generating rationales from the best performing 25-shot prompt (with model-generated rationales) on the MATH training set and using the problems which were not solved in first iteration – seem to further improve many-shot performance.
@@ A.3 Additional Results for Reinforced and Unsupervised ICL
A.4 Linear Classification: Data Generation
@@
A.5 Example Prompts
Figure A.20: An example 1-shot PDDL (Ghallab et al., 1998) prompt, with a test example for the Logistics domain in §$3.$ Within a city, the locations are directly linked, allowing trucks to travel between any two of these locations. Similarly, cities are directly connected to each other allowing airplanes to travel between any two cities. Each city is equipped with one truck and has a designated location that functions as an airport
Figure A.21: Zero-shot prompt for GPQA.
Figure A.22: An example 2-shot code correctness verifier prompt on GSM8K. Here, the solutions are generated using the Gemini 1.0 Pro (Gemini Team, 2023) model. See results in §3.4.
Figure A.23: 4-Shot Inner Monologue prompt used for MATH and GSM8K.
Figure A.24: Example prompt with 8 shots per class for the linear classification in 16 dimensions, discussed in §5.2. Here, we use semantically-unrelated labels (‘Foo’ and ‘Bar’) following Wei et al. (2023).
A.6 Prompts for Unsupervised ICL
@@
A.7 Training GPT-2 from scratch on the sequential parity task
[…]

### v2

@@ 4.6 Long-context scaling laws may not predict ICL performance
Figure 16: Negative Log-Likelihood (NLL) as a function of number of shots. We plot NLL on ground truth test set solutions for GPQA, MATH and GSM8K. For GPQA and MATH, questions for Reinforced ICL and Unsupervised ICL comes from the training splits of those datasets. We study GSM8K in the transfer setting, i.e. questions for Reinforced and Unsupervised ICL come from MATH. The absolute NLL for ICL and Reinforced ICL are not directly comparable to Unsupervised ICL, since they use different prompt formats.
@@
Prior works [68, 2, 27] have found that the negative log-likelihood (NLL) for ground-truth test outputs decreases predictably as the context length increases. We confirm this finding for GPQA, Hendrycks MATH and GSM8K with many-shot ICL, and report our results in Figure 16. However, we note that NLL trends are not a strong predictor for downstream task performance. For example, the success rate for both MATH and GPQA with ICL decreases after 125 shots (Figure 7,8), but we do not observe a corresponding increase in the NLL in Figure 16.
We also plot NLL curves for Reinforced and Unsupervised ICL, and find them to generally have a smaller slope when compared to supervised ICL. Interestingly, NLL curves for ICL with ground-truth outputs is lower than with model-generated outputs, even though the latter often performs better. In the GSM8K transfer setting (using MATH problems and solutions to score GSM8K solutions), the change in NLL is close to nil. However, this doesn’t reflect transfer performance on GSM8K, which continues to improve with more examples (Figure 7).
@@
Overall, our results demonstrate that NLL is not a reliable proxy when attempting to predict ICL performance for problem-solving domains. This makes intuitive sense: for any given problem, there are a large number of potentially correct CoT solutions that the model can generate, and calculating the log-likelihood on only one such solution may not provide a clear picture for overall model capability. We also explore computing NLL on a diverse set of model-generated outputs on MATH, and our findings are presented in §A.7.
5 Related Work
Scaling in-context learning
Brown et al. [8] reported improved performance as you increase the number of examples (up to 64) for in-context learning in LLMs , and later works corroborated this finding [39]. However, very few works have explored using a large number of examples (1000 or above) in the prompt. This is likely due to the fact the context lengths in large language models have been quite limited until recently [16, 3]. One closely related work to ours is from Li et al. [31], who scale the number of examples for in-context learning to 2000. However, Li et al. [31] use a custom model architecture [74] to achieve long context lengths, and only evaluate models of up to 1.3B parameters, which is several orders of magnitude smaller than state-of-the-art language models, and are ineffective for complex tasks, such as GPQA [52].
Concurrently to our work, Anil et al. [2] used many-shot prompting (upto 256 shots) to jailbreak language models. In our work, we focus on a much wider range of tasks, use a lot more examples (up to 8192 shots) and use models with much longer context lengths (up to 1M tokens). Also, we explore mitigations for needing many human-generated examples with many-shot ICL. Furthermore, while Anil et al. [2] use many-shot learning to override preferences learned during RLHF phase to elicit the biases stemming from pretraining, our results in §4.1 demonstrate that we can also override pre-training biases themselves. Bertsch et al. [6] also concurrently shows benefits of scaling up in-context learning to many demonstrations on several classification datasets with up to 151 labels, albeit also using smaller context windows of up to 80k tokens (using Llama2-80k [13]).
Long-context scaling laws
Prior works [68, 2, 27, 16] have reported smaller next-token prediction loss with longer contexts, which Jeon et al. [25] also show using theoretical analysis. Our findings confirm this trend for even longer context lengths, but our analysis reveals some of the limitations of using next-token prediction loss as a metric for evaluating long-context performance, as next-token prediction loss continues to go down even as overall performance plateaus.
Learning from self-generated data
Numerous recent works [19, 70, 55] propose fine-tuning language models on self-generated data to improve performance. Their approach consists of (1) generate samples from the model and filter them using binary feedback, (2) fine-tune the model on these samples, and (3) repeat this process a few times. In this work, we extend this idea to in-context learning, and study the efficacy of Reinforced ICL in reasoning and problem-solving domains.
Self-generated data and in-context learning
Kim et al. [29] propose using self-generated data for few-shot ICL on classification problems, where they generate demonstrations using the LLM conditioned on the test input for each possible class label, and including these demonstrations in the context when performing the final prediction. Li et al. [32] extend this approach to reasoning and language understanding tasks, where they also generate demonstrations conditioned on the test input. Consistent with our findings, these works show that model-generated demonstrations can outperform human-generated demonstrations in the few-shot regime. Another related approach is AutoCoT [73] that uses a zero-shot CoT prompt to produce model-generated demonstrations for few-shot ICL. To do so, AutoCoT samples diverse questions one-by-one based on embedding-based clustering followed by heuristics-based post-processing for selecting demonstrations.
Different from above approaches, Reinforced ICL generates demonstrations using the same procedure as Singh et al. [55], does not require clustering, post-processing heuristics, or access to the test inputs for generating demonstrations, and can be applied to any problem for which we can obtain reliable reward signals. Moreover, our work mainly focuses on the utility of randomly-sampled model-generated demonstrations for many-shot ICL.
Learning Input-Output Relationships with ICL
Numerous works [41, 30, 69, 36] have investigated whether LLMs truly learn input-output relationships during in-context learning. Min et al. [41] found that replacing the ground truth labels in in-context examples with random labels barely effected final performance. Further investigations by Yoo et al. [69] and Kossen et al. [30] found that this finding does not necessarily hold across tasks and model sizes. In particular, Kossen et al. [30], Lin and Lee [36] showed that LLMs can indeed learn input-output relationships via in-context learning, but require more examples in order to do so well. In our work, we extrapolate the trend found in those works to much longer context lengths, showing that pre-training biases can be mostly overcome given enough training examples.
Learning Mathematical Functions with LLMs
Several prior works investigate whether mathematical functions can be learned with transformers [14, 72, 67, 7]. All these works train transformers specifically to perform in-context learning for such functions. In contrast, we demonstrate that many-shot ICL can learn high-dimensional functions even with pre-trained LLMs. Concurrent to our work, Vacareanu et al. [58] demonstrate that pretrained LLMs are able to perform regression tasks, with performance rivaling that of traditional supervised methods with 500 in-context examples. Our work complement their findings to other synthetic tasks with a much larger number of in-context examples. Dinh et al. [11] fine-tuned GPT-3 on synthetic classification tasks and observed similarities in the decision boundaries learned by the fine-tuned model and kNNs. Our results in Figure 11 show that many-shot ICL also performs comparably to kNNs on high-dimensional classification tasks.
Comparing ICL with fine-tuning
Contrary to task-specific fine-tuning, ICL does not require optimizing any model weights, allowing LLMs to perform a variety of tasks at inference. As such, several prior works compare fine-tuning with ICL but in few-shot regime. Liu et al. [37] proposed a parameter-efficient few-shot fine-tuning (FT) approach for T0 that outperforms few-shot ICL with GPT-3. However, Awadalla et al. [5] argue that few-shot ICL is more robust to distribution shifts than fine-tuning for question answering tasks. Similarly, Asai et al. [4] show better transfer with ICL compared to fine-tuning on some tasks. Mosbach et al. [42] fairly compare ICL with FT by using the same model for both approaches and show that full fine-tuning (FT) generally outperforms ICL in the few-shot regime with 16 examples. More recently, Lin et al. [34] show that few-shot ICL can outperform fine-tuning based approaches for aligning LLMs.
Complementary to prior works, we compare full fine-tuning with many-shot ICL with the same number of examples for low-resource translation. Notably, we find that many-shot ICL performs comparably to FT. Aligned with our findings, Bertsch et al. [6] concurrently show that many-shot ICL generally outperforms parameter-efficient fine-tuning (LoRA) on classification tasks. Overall, many-shot ICL and FT can exhibit comparable behaviors, which we leave for further investigation.
Exemplar vs. Rule-based ICL generalization
Chan et al. [9] indicate that ICL tends to generalize in a more exemplar-based way, compared to rule-based generalization during in-weights learning. Using a clever experiment with blocked attention, Bertsch et al. [6] also argue that the benefits of many in-context demonstrations arise from having access to more similar examples. While our results on in-context linear classification agree with this conclusion, our sequential parity results seem to contradict it. Strikingly, sequential parity was the task on which we saw the most improvement, whereas it should be a task that benefits least from seeing similar examples – after all, the nearest neighbor is always going to give the wrong answer (off by 1 bit). Chan et al. [9] do show that a transformer’s inductive biases towards exemplar-based generalization can be shifted both by the training data and the model size, with larger models being less exemplar-based – perhaps this explains the contradictory findings, given that our work used a larger and much more capable model, though this remains an open question.
6 Discussion, Limitations and Future Work
@@
We found significant gains in performance when going from few-shot to many-shot ICL on a wide range of tasks, including translation, summarization, planning, reward modeling, mathematical problem solving, question-answering, algorithmic reasoning, and sentiment analysis. To overcome the challenges of obtaining a large number of high-quality human-written rationales for many-shot ICL, we introduced two regimes: Reinforced ICL and Unsupervised ICL. Moreover, we demonstrate that, unlike few-shot ICL, many-shot ICL is effective at overriding pretraining biases, can learn high-dimensional functions with numerical inputs, and performs comparably to SFT.
@@
One limitation of our work is that it mainly evaluates many-shot ICL with Gemini 1.5 Pro. That said, concurrent works [2, 6] as well as our preliminary results with GPT-4-Turbo and Claude-3-Opus (§4.5) indicate that other LLMs can also benefit from many-shot ICL. Future work should focus on evaluating the many-shot abilities of a wide range of long context models, as they become available. Furthermore, many-shot performance can likely serve as a valuable metric for evaluating the quality of long-context models, going beyond the prevalent needle-in-a-haystack test [26].
@@
Another limitation of our work is that we don’t completely understand why performance can sometimes degrades with more examples in the prompt (for example, for MATH). Our analysis found that negative log-likelihood trends are insufficient to explain this degradation, and future work should investigate new directions to shed light on the matter and improving many-shot ICL capabilities. Overall, we hope that this work lays a foundation for understanding and optimizing the use of long-context models for ICL, opening up a new frontier of LLM capabilities.
@@ Acknowledgements
We would like to thank Gheorghe Comanici for reviewing an early draft of this work. We are also grateful to Doina Precup, Aviral Kumar, Dale Schuurmans, Ankit Anand, Ross Goroshin, Urvashi Singh, Jannik Kossen, Charline Le Lan, and Daniel Toyoma for helpful discussions.
@@
RA initiated and led the project, ran majority of the many-shot experiments and analysis, came up with reinforced ICL, on-boarded collaborators, wrote the initial draft. AS contributed initial infra for experiments on MATH and GSM8K, ran BBH experiments, co-led the fine-tuning experiments, conducted NLL analysis on problem-solving tasks, and wrote several sections.
LZ contributed results for in-context verifier. BB contributed the planning logistics task. LR led the fine-tuning experiments. BZ contributed the many-shot results for GPT-4 and Claude-3. AA helped with GPQA, SC contributed the baseline for parity task and both helped edit the paper. AF and HL provided feedback on an early draft. HL also suggested the unsupervised ICL experiments. Others were involved in project discussions and minor edits to the paper.
@@ Appendix A Appendix
A.1 Example Prompts
Figure A.3: An example 1-shot PDDL [17] prompt, with a test example for the Logistics domain in §2.3. Within a city, the locations are directly linked, allowing trucks to travel between any two of these locations. Similarly, cities are directly connected to each other allowing airplanes to travel between any two cities. Each city is equipped with one truck and has a designated location that functions as an airport
Figure A.4: Zero-shot prompt for GPQA.
Figure A.5: An example 2-shot code correctness verifier prompt on GSM8K. Here, the solutions are generated using the Gemini 1.0 Pro [15] model. See results in §2.4.
Figure A.6: 4-Shot Inner Monologue prompt used for MATH and GSM8K.
Figure A.7: Example prompt with 8 shots per class for the linear classification in 16 dimensions, discussed in §4.2. Here, we use semantically-unrelated labels (‘Foo’ and ‘Bar’) following Wei et al. [63].
A.2 Prompts for Unsupervised ICL
[…]
A.3 Unsupervised ICL on Machine Translation
@@
A.4 Reinforced ICL: Data-collection Prompt Sensitivity and Iteration 2
Figure A.12: Reinforced ICL Hendrycks MATH. We find the performance of model-generated rationales with 4-shot Minerva prompt is generally better or comparable to the ones generated by 4-shot InnerMono. MATH prompt. Furthermore, another iteration of Reinforced ICL – generating rationales from the best performing 25-shot prompt (with model-generated rationales) on the MATH training set and using the problems which were not solved in first iteration – seem to further improve many-shot performance.
A.5 Linear Classification: Data Generation
For each classification dataset, we randomly sample another $N$-dimensional vector as the decision boundary and a decision threshold. We then provide $K$ N-dimensional points above this threshold and $K$ points below that same threshold as in-context exemplars, and the model must determine whether unseen N-dimensional points are above or below the threshold (we do not tell the model the equation or the threshold). We provide the python code for date generation below.
@@ A.5 Linear Classification: Data Generation
A.6 Training GPT-2 from scratch on the sequential parity task
@@
A.7 Negative Log-Likelihood on Model-Generated Data
A.8 Many-Shot ICL: Hallucinations on Summarization
Here is an example on XSum showing summaries with halluncinated date and time when increasing the number of shots for many-shot ICL, discussed in §2.2.
50-shot: 12 September 2016 Last updated at 14:32 GMT
One of Russia’s last independent pollsters has been designated a "foreign agent" by the justice ministry.
125-shot: 12 September 2016 Last updated at 14:43 GMT
One of Russia’s last independent pollsters has been branded a "foreign agent" by the justice ministry.
250-shot: 2016-09-06 Last updated at 14:33 GMT
Russia’s only independent pollster has been branded a "foreign agent" by the justice ministry.
500-shot: 12 September 2016 Last updated at 14:22 GMT
One of Russia’s last independent pollsters has been branded a "foreign agent" by the government.